The vector valued function r(t)=⟨t^3+4,e^t,cos(t)⟩

is
Select one:
a. continuous for all t∈R

b. only continuous for t∈[0,∞)


c. continuous for all t
except 0
d. only continuous at t=Ï€


e. only continuous for t∈[1,∞)
1
Expert's answer
2020-05-19T19:10:34-0400

Theorem. "r(t)=\big(f(t), g(t), h(t)\big)" is continuous at "t=a" if and only if all of "f, g, h" are continuous at "t=a."

"f(t)=t^3+4" is continuous everywhere,

"g(t)=e^t" is continuous everywhere,

"h(t)=\cos(t)" is continuous everywhere.

The vector valued function "r(t)=\big(t^3+4, e^t, \cos(t)\big)" is

a. continuous for all t∈R